The first day at a baseball game you purchase 3 colas and 2 hotdogs worth $5.75.
The next day you purchase 2 colas and 5 hotdogs worth $7.50. Find the cost of one cola and one hotdog.

Answer by mananth(16949)

let the cost of cola be $x
the cost of burger be $y
First day
3x+2y= 5.75
Second day
2x+5y=7.50
.
2(3x+2y)-3(2x+5y)=2*5.75 - 3*7.50
6x+4y-6x-15y=11.50-22.50
-11y=-11
y= 1
y=$1.00 the cost of a burger
Plug the value of y
3x+2=5.75
3x=3.75
x=$1.25 cost of a cola